OpenStudy (anonymous):

Find the equation of a straight line that passes through the point (2,4) and is parallel to the line 2x+3y-8=0

OpenStudy (anonymous):

first, find the slope of the line: clue: Slope (m) of the equation of a straight line is same as the slope of 2x+3y-8=0, since they are parallel. second, plug the slope in y=mx+b (slope intercept form) third, using the given point (2,4) find the y-intercept of the equation (b) lastly, rewrite the slope(m) and y-int(b) into y=mx+b (slope intercept form).

OpenStudy (anonymous):

Okay I'll try that thanks!!

OpenStudy (anonymous):

yw(^_^)